OpenStudy (anonymous):

. need help (d) William’s wife says the molding for the window will take almost 50% of the total molding they need. William says it is closer to of the molding. Approximately what percent of the total molding will be used for the window and who is closer to the correct answer? Explain your answer.

OpenStudy (anonymous):

ok, so the total molding they need is 65 1/8 , and

OpenStudy (anonymous):

the amount of molding they need for the window is 25 7/8

zepdrix (zepdrix):

Williams says it is closer to ____ of the molding. It looks like that didn't copy over correctly.

OpenStudy (anonymous):

oh sorry i'll add that in

OpenStudy (anonymous):

1/3

zepdrix (zepdrix):

50% is half of the total amount, we can write this proportion as 1/2. So we're trying to determine if the amount of framing needed is closer to 1/2 or 1/3 of the total given. Err not framing, molding, whatever :)

zepdrix (zepdrix):

So this is a division problem. You divide the `amount needed` by the `total amount` and that will tell you what proportion of the total you are using. \(\large\rm 65\frac{1}{8}\div25\frac{7}{8}\)

zepdrix (zepdrix):

Woops* I wrote that division backwards >.< my bad

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):

We want to divide the amount used by the total, \(\large\rm 25\frac{7}{8}\div65\frac{1}{8}\)

zepdrix (zepdrix):

So maybe we should rewrite these as improper fractions? Mixed numbers are no fun.

zepdrix (zepdrix):

Do you know how to do that? :o

zepdrix (zepdrix):

\(\large\rm 25\frac78=\dfrac{?}{8}\)

OpenStudy (anonymous):

Yes, they're 521/8 and 207/8, right?

zepdrix (zepdrix):

good good good :)

zepdrix (zepdrix):

\[\large\rm \frac{207}{8}\div\frac{521}{8}\]Have you ever heard of `Keep Change Flip`? It's a popular device used to help remind students how to deal with division of fractions.

OpenStudy (anonymous):

Yup! I know that. Flip the 2nd fraction and change it to multiplication

zepdrix (zepdrix):

Mmm k good good good.

zepdrix (zepdrix):

\[\large\rm \frac{207}{8}\div\frac{521}{8}\quad=\frac{207}{8}\times\frac{8}{521}\]

OpenStudy (anonymous):

then cross cancel?

zepdrix (zepdrix):

Ooo yes, seems like a good idea!

zepdrix (zepdrix):

\[\large\rm \frac{207}{521}\]

OpenStudy (anonymous):

so turn that into a mixed fraction?

zepdrix (zepdrix):

Hmm this is annoying, this one is going to be right in the middle between 1/3 and 1/2 lol. Will require some extra work to determine which is closer.

zepdrix (zepdrix):

We have no `whole portion` in this fraction :) So no mixed number

zepdrix (zepdrix):

So umm

OpenStudy (anonymous):

how did you determine whether it's close or not?

zepdrix (zepdrix):

I cheated, I used a calculator. Here is an idea... This is going to be require some big numbers :) but... I guess you could look for a common denominator between these three fractions,\[\large\rm \frac{1}{2}\qquad \qquad\frac{1}{3}\qquad \qquad\frac{207}{521}\]and then simply compare the numerators, see which two are closest together. Make sense? :o

OpenStudy (anonymous):

How did you cheat your way through in finding the solution (using the calculator)

zepdrix (zepdrix):

I know that 1/2 = 0.5 and that 1/3 = 0.333 I typed 207/521 into the calculator and got = 0.39731286

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so other than that, should I find the common denominator and then compare the numerators to tell whose closer?

zepdrix (zepdrix):

521 is a prime number, so your common denominator would need to be 2x3x521

zepdrix (zepdrix):

Ya that might be a fun way to finish it up ^^

OpenStudy (anonymous):

so 1/3??

zepdrix (zepdrix):

\[\large\rm \frac{1}{2}\qquad \qquad\frac{1}{3}\qquad \qquad\frac{207}{521}\]So your numerators end up being 3*521, 2*521 and we want to know which is closer to 2*3*207. Yup! Looks like it was the 1/3, ya? \c:/

OpenStudy (anonymous):

ok, thank you!